Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-3x + 4y = -19.5
-3x + y = -10.5

Answer by mananth(16946)

-3 x + 4 y = -19.5 .............1
-3 x + 1 y = -10.5 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -4
-3 x 4 y = -19.5
12 x -4 y = 42
Add the two equations
9 x = 22.5
/ 9
x = 2.50 OR 2 1/ 2
plug value of x in (1)
-3 x + 4 y = -19.5
-7.5 + 4 y = -19.5
4 y = -19.5 + 7.5
4 y = -12
y = -3
(5/2,-3)